OpenStudy (anonymous):

find the area of a circle with a circumference of 10*pie*sqrt 3

OpenStudy (anonymous):

once again the answer is 300*pie but I have no idea how to get it

OpenStudy (kc_kennylau):

Find its radius first

OpenStudy (kc_kennylau):

C=2*pi*r

OpenStudy (anonymous):

10*pie*sqrt3= 2*pie*r so radius = 5*sqrt3

OpenStudy (kc_kennylau):

Then find its area by A=pi*r*r

OpenStudy (anonymous):

75*pi right?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

but it was on a test and the answer is 300*pi

OpenStudy (kc_kennylau):

Then the answer is wrong

OpenStudy (ipwnbunnies):

*applause* I love technology.

OpenStudy (ipwnbunnies):

Did you type your question out wrong?

OpenStudy (anonymous):

idk how do you attach a windows document

OpenStudy (ipwnbunnies):

'Attach File' near the bottom of the question.

OpenStudy (anonymous):

I cant find it in my documents but the question is perfectly typed

OpenStudy (ipwnbunnies):

Your assignment is a lie and you should smite the creator.

OpenStudy (anonymous):

hes a 65 year old man who looks like walt from breaking bad

OpenStudy (ipwnbunnies):

At least he's got that going for him lol.

OpenStudy (anonymous):

questions 6 and 7 I was having trouble with

OpenStudy (ipwnbunnies):

Dude, number 7 says the area of a circle with circumference 10pi....

OpenStudy (ipwnbunnies):

Oh, sqrt(3) just popped out of nowhere on that question. I love technology.

OpenStudy (anonymous):

ill go to class early on Monday and show him the error of his ways

OpenStudy (ipwnbunnies):

Show him, tell him to say your name. He'll get it. 8l

OpenStudy (anonymous):

this website is awesome